OpenStudy (anonymous):

A can of peanuts has a circular base and top, and it is cylindrical in shape. The volume is 500 cm3. If the can is to require the least amount of material, what must its dimensions be? Give your answer correct to one decimal place.

OpenStudy (anonymous):

what do you get for r ?

OpenStudy (accessdenied):

What have you considered on this problem so far?

OpenStudy (accessdenied):

The volume is given. If you know the formula for volume, you can set up one relationship between the variables of radius and height. Then, to minimize the amount of materials, the only relevant part of the can using material is the surface itself, so we are minimizing surface area essentially.

OpenStudy (anonymous):

i used the area formula after ? is that wrong ?

OpenStudy (accessdenied):

What area formula did you use?

OpenStudy (anonymous):

2pirh+2pir^2

OpenStudy (accessdenied):

That looks good to me. Then we just need to use the first volume equation to eliminate one variable.

OpenStudy (anonymous):

so this is where i am stuck r^3=250/pi

OpenStudy (accessdenied):

\( V = \pi r^2 h \) This is the volume of a cylinder. \( 500 = \pi r^2 h \) We need to solve for one of these variables to use in surface area, which allows us to minimize surface area in terms of one dimension.

OpenStudy (anonymous):

yep i did this h=500pir^2

OpenStudy (anonymous):

500/pir^2 i mean

OpenStudy (accessdenied):

So you substituted that into surface area: \(S = 2 \pi r^2 + 2 \pi r h \) \( h = \dfrac{500}{\pi r^2} \) \(S = 2 \pi r^2 + 2 \pi r \left( \dfrac{500}{\pi r^2} \right) \)

OpenStudy (accessdenied):

With this equation, we can use optimization to find the value for r that minimizes the surface area.

OpenStudy (anonymous):

yea so i was stuck at r^3=250/pi

OpenStudy (accessdenied):

\( S = 2 \pi r^2 + 2\cancel{\pi r } \dfrac{500}{\cancel{\pi r} r} \) \( S = 2 \pi r^2 + \dfrac{1000}{r} \) \( S' = 4 \pi r - \dfrac{1000}{r^2} = 0 \) \( 4 \pi r = \dfrac{1000}{r^2} \) \( r^3 = \dfrac{1000}{4 \pi } \) You would take the cube root from here, or power to 1/3. \( \sqrt[3]{r^3} = \sqrt[3]{\dfrac{250}{\pi}} \) That makes sense?

OpenStudy (anonymous):

okey yea then next step ??

OpenStudy (anonymous):

i know we sub it in

OpenStudy (accessdenied):

That gives us the radius. Substituting it into the volume formula, then: \( 500 / \left[ \pi \left( \sqrt[3]{\dfrac{250}{\pi}} \right) ^2\right] = h \)

OpenStudy (accessdenied):

I presume "dimensions of the can" implies diameter and height. So we could double the radius for diameter.

OpenStudy (anonymous):

what do u get after that step ??

OpenStudy (anonymous):

do u get 12.599 for h ?

OpenStudy (accessdenied):

I obtained h = 8.602 according to calculator.

OpenStudy (anonymous):

wait what did u do ?

OpenStudy (accessdenied):

\( 500 / \left[ \pi \left( \sqrt[3]{\dfrac{250}{\pi}} \right) ^2\right] = h \) This has to be done on a calculator, being careful of the order in which items are attended to. 500/(pi*(250/pi)^(2/3))

OpenStudy (anonymous):

i canclced the pi thats what i did wrong i think

OpenStudy (accessdenied):

Yeah, those two pi's don't have the same degree since one is raised to the 2/3 power. The main simplification would be to cancel 2/3 from both leaving \( \pi^{1/3} \), but it is usually easier to just enter it all in a calculator to avoid those mess-ups.